# Quality control in sublinear time: a case study via random graphs

### v1

@@ 5.1.2 Inductive Lemma
From this, Section 5.1.2 implies that with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, the uniformly sampled multiset $S$ satisfies $C_{H^{\prime}}(S)\in(1\pm\frac{\varepsilon}{4})S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$ for all $H^{\prime}\subseteq H$. Therefore, the algorithm accepts $G\sim\mathcal{D}_{n}$ with high constant probability.
@@ A.1 Query complexity upper bound
There is an algorithm that uses $O\left(\sqrt{n/d}\right)$ degree queries and time and estimates the average degree of an input graph $G$ up to a multiplicative factor of $3$, with high constant probability, as long as $G$ has average degree at least $d$.
@@
Consider an input graph $G$ with $e(G)$ edges, and a parameter $\beta$. There exists an algorithm (call it Algorithm B) that performs expected $O(\log(n)/p)$ queries and has expected runtime $O(\log(n)/p)$ with the following guarantees. If $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability. If $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>10n^{2}p$, the algorithm rejects with high constant probability.
@@
Let $\rho_{e}(G)=e(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[e(G^{\prime})\right]$. While the focus in this section is on triangles, a modification of this corollary already implies a $(\mathcal{G}_{n,p},\rho_{e})$-quality control algorithm for comparing the number of edges to $\mathcal{G}_{n,p}$, with $\widetilde{O}(\sqrt{1/p})\cdot\text{poly}(1/\varepsilon)$ adjacency list queries and runtime. Particularly, [GR08] provides stronger guarantees than [Fei06], giving a $\widehat{O}(n^{2}/e(G))^{1/2}\cdot\text{poly}(1/\varepsilon)$-time algorithm that outputs a $(1\pm\varepsilon)$-approximation of the average degree in a graph $G$. We can use this to prove a more refined version of Section A.1. For quality control, the completeness case follows from the concentration of the number of edges for $G^{\prime}\sim\mathcal{G}_{n,p}$. Soundness follows from rejecting graphs whose edge count is out of $(1\pm\varepsilon)\binom{n}{2}p$.
@@
If the first step does not reject, we run the second step, i.e. the procedure from [Fei06]. If the input graph has $e(G)\geq\frac{1}{5}n^{2}p$, by Lemma A.4 this step takes $O(\sqrt{1/p})$ time. Otherwise, the algorithm can take $O(n^{2})$ time; this will happen with probability at most $1/n^{2}$. The algorithm has expected $O(\sqrt{1/p})$ queries and has expected runtime $O(\sqrt{1/p})$. This procedure has the guarantee that if $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability, and if $e(G)>10n^{2}p$, the algorithm rejects with high probability. ∎
@@
When $\mathrm{arb}(G)\leq\alpha^{\prime}$, if $C_{3}(G)\geq t$, then with high constant probability, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
@@
Completeness: We first reason about completeness, i.e., that Algorithm 1 accepts $G\sim\mathcal{G}_{n,p}$ with high probability (both over the distribution over graphs and the algorithm’s randomness). By Section 2.3 and Section 2.3, with high constant probability, since $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$, $G\sim\mathcal{G}_{n,p}$ satisfies $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$.
@@
The algorithm above only used the following properties of $\mathcal{G}_{n,p}$: with high constant probability, the arboricity is bounded, and the number of edges and triangles are each concentrated around their expected values. Therefore, the algorithm and analysis extend to any distribution over graphs with these properties. Indeed, this is a recurring theme in our results: we utilize minimal structural properties about $\mathcal{G}_{n,p}$ and the results extend nicely to other distributions over graphs that satisfy these properties.
@@ B.2 Quality control algorithm and analysis
Completeness: Suppose the input graph $G$ is drawn from $\mathcal{G}_{n,p}$. Then, by the completeness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, each call of Algorithm Induced-Motif-Quality-Efficient in Step 1 accepts $G$ with probability at least $1-\frac{1}{3\cdot M(kd)}$. (Observe that Algorithms 2 and 3 were analyzed with a different constant success probability, but since $M(k,d)$ is constant with respect to $n,p$, we can replace it with $1-\frac{1}{3\cdot M(k,d)}$ in the analysis.) By a union bound over the $M(k,d)$ possible induced subgraphs over at most $2k^{d}$ vertices, Step 2 then accepts $G\sim\mathcal{G}_{n,p}$ with probability at least $2/3$.
@@ C.1 From average-case complexity to quality control and back
Works on average-case algorithms typically define a “good” property of inputs with the following guarantees. With probability at least $1-1/t_{2}$, an input drawn from $\mathcal{D}$ possesses the good property, and thus there exists an algorithm $A$ with worst-case runtime $t_{1}$ for this input. This “good” property is possible to certify/verify. On the other hand, for worst-case inputs, the algorithm $A$ has worst-case runtime $t_{2}$. In such a set-up, the average-case complexity of the algorithm $A$ is:
@@
We now argue about the runtime. First, with probability at least 11/T1-1/T, $x\sim\mathcal{D}$ satisfies $\rho(x)\in(1\pm\varepsilon)$, and so with probability at least $(1-1/T)\cdot 2/3\geq 1-1/t_{4}$ (over the input distribution), $x\sim\mathcal{D}$ is accepted by Algorithm $C$. In this case, the algorithm takes time $t_{3}$. If $x$ is not accepted by Algorithm $C$, then Algorithm $E$ is called, and will take time $t_{4}$. Therefore, the average-case runtime of the algorithm is:
@@ D.1 Erdős–Rényi graphs
Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
@@
Since $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, this gives us:
$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{\varepsilon^{2}}{10}.$
@@
$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}\leq\frac{1}{10}.$

### v2

@@ 5.1.2 Inductive Lemma
From this, Section 5.1.2 implies that with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)=1-o(1)$, the uniformly sampled multiset $S$ satisfies $C_{H^{\prime}}(S)\in(1\pm\frac{\varepsilon}{4})S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$ for all $H^{\prime}\subseteq H$. Therefore, the algorithm accepts $G\sim\mathcal{D}_{n}$ with high probability.
@@ A.1 Query complexity upper bound
There is an algorithm that uses $\widetilde{O}\left(\sqrt{n/d}\right)$ degree queries and time and estimates the average degree of an input graph $G$ up to a multiplicative factor of $3$, with probability 1o(1)1-o(1), as long as $G$ has average degree at least $d$.
@@
Consider an input graph $G$ with $e(G)$ edges, and a parameter $\beta$. There exists an algorithm (call it Algorithm B) that performs expected $O(\log(n)/p)$ queries and has expected runtime $O(\log(n)/p)$ with the following guarantees. If $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high probability. If $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>10n^{2}p$, the algorithm rejects with high probability.
@@
Let $\rho_{e}(G)=e(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[e(G^{\prime})\right]$. While the focus in this section is on triangles, a modification of this corollary already implies a $(\mathcal{G}_{n,p},\rho_{e})$-quality control algorithm for comparing the number of edges to $\mathcal{G}_{n,p}$, with $\widetilde{O}(\sqrt{1/p})\cdot\text{poly}(1/\varepsilon)$ adjacency list queries and runtime. Particularly, [GR08] provides stronger guarantees than [Fei06], giving a $\widetilde{O}(n^{2}/e(G))^{1/2}\cdot\text{poly}(1/\varepsilon)$-time algorithm that outputs a $(1\pm\varepsilon)$-approximation of the average degree in a graph $G$. We can use this to prove a more refined version of Section A.1. For quality control, the completeness case follows from the concentration of the number of edges for $G^{\prime}\sim\mathcal{G}_{n,p}$. Soundness follows from rejecting graphs whose edge count is out of $(1\pm\varepsilon)\binom{n}{2}p$.
@@
If the first step does not reject, we run the second step, i.e. the procedure from [Fei06]. If the input graph has $e(G)\geq\frac{1}{5}n^{2}p$, by Lemma A.4 this step takes $O(\sqrt{1/p})$ time. Otherwise, the algorithm can take $O(n^{2})$ time; this will happen with probability at most $1/n^{2}$. The algorithm has expected $O(\sqrt{1/p})$ queries and has expected runtime $O(\sqrt{1/p})$. This procedure has the guarantee that if $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high probability, and if $e(G)>10n^{2}p$, the algorithm rejects with high probability. ∎
@@
When $\mathrm{arb}(G)\leq\alpha^{\prime}$, if $C_{3}(G)\geq t$, then with high probability, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
@@
Completeness: We first reason about completeness, i.e., that Algorithm 1 accepts $G\sim\mathcal{G}_{n,p}$ with high probability (both over the distribution over graphs and the algorithm’s randomness). By Section 2.3 and Section 2.3, with high probability, since $p=\omega\left(\varepsilon^{-2}\cdot 1/n\right)$, $G\sim\mathcal{G}_{n,p}$ satisfies $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$.
@@
The algorithm above only used the following properties of $\mathcal{G}_{n,p}$: with high probability, the arboricity is bounded, and the number of edges and triangles are each concentrated around their expected values. Therefore, the algorithm and analysis extend to any distribution over graphs with these properties. Indeed, this is a recurring theme in our results: we utilize minimal structural properties about $\mathcal{G}_{n,p}$ and the results extend nicely to other distributions over graphs that satisfy these properties.
@@ B.2 Quality control algorithm and analysis
Completeness: Suppose the input graph $G$ is drawn from $\mathcal{G}_{n,p}$. Then, by the completeness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, each call of Algorithm Induced-Motif-Quality-Efficient in Step 1 accepts $G$ with probability at least 1o(1)1-o(1). By a union bound over the $M(k,d)$ (a constant) possible induced subgraphs over at most $2k^{d}$ vertices, Step 2 then accepts $G\sim\mathcal{G}_{n,p}$ with probability 1o(1)1-o(1).
@@ C.1 From average-case complexity to quality control and back
Works on average-case algorithms typically define a “good” property of inputs with the following guarantees. With probability at least $1-1/t_{2}$, for $t_{2}=\omega(1)$, an input drawn from $\mathcal{D}$ possesses the good property, and thus there exists an algorithm $A$ with worst-case runtime $t_{1}$ for this input. This “good” property is possible to certify/verify. On the other hand, for worst-case inputs, the algorithm $A$ has worst-case runtime $t_{2}$. In such a set-up, the average-case complexity of the algorithm $A$ is:
@@
We now argue about the runtime. First, with probability at least 11/T1-1/T, $x\sim\mathcal{D}$ satisfies $\rho(x)\in(1\pm\varepsilon)$, and so with probability at least $(1-1/T)\cdot(1-o(1))\geq 1-1/t_{4}$ (over the input distribution), $x\sim\mathcal{D}$ is accepted by Algorithm $C$. In this case, the algorithm takes time $t_{3}$. If $x$ is not accepted by Algorithm $C$, then Algorithm $E$ is called, and will take time $t_{4}$. Therefore, the average-case runtime of the algorithm is:
@@ D.1 Erdős–Rényi graphs
Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $n\geq\frac{\omega_{1/p}(1)}{\varepsilon^{2}}\cdot p^{-m_{H}}$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
For $G\sim\mathcal{G}_{n,p}$, with probability at least $1-o_{1/p}(1)$, the number of copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
@@
Since $n\geq\frac{\omega_{1/p}(1)}{\varepsilon^{2}}\cdot p^{-m_{H}}$, this gives us:
$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{\varepsilon^{2}}{\omega(1)}=\mathbb{E}\left[X_{G}\right]^{2}\cdot\varepsilon^{2}\cdot o_{1/p}(1).$
@@
$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}=o_{1/p}(1).$